OpenStudy (anonymous):

Find the domain of this rational function R(x) = 5/x^2-7x I am so confused

OpenStudy (anonymous):

tell me what does domain mean

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

domain is the all the numbers that can be put in the equation and gives you a real number

OpenStudy (anonymous):

\[\frac{ 5 }{x^2-7x }=\frac{ 5 }{ x(x-7) }\]

OpenStudy (anonymous):

a number can not be divided by a 0 so what values give the bottom 0

OpenStudy (anonymous):

Ooooh okay, that x after the 7 was throwing me off

OpenStudy (anonymous):

I know what domain is, I just can't explain it. Like, I know what I was supposed to do with the equation I just didn't know how to. Thanks!

OpenStudy (anonymous):

np glad i could help :D